# TESSERA: Temporal Embeddings of Surface Spectra for Earth Representation and Analysis

### v2

@@ 2.2 Canopy height estimation
Despite substantial progress in global canopy height mapping, current approaches exhibit limitations that restrict their reliability and applicability. GEDI, a spaceborne LiDAR system, can provide reasonably accurate canopy height measurements to train models to generate CHMs from optical and radar data; however, it does not provide wall-to-wall maps, and does not sample high latitudes [41], and produces unreliable estimates in mountainous terrain [52]. Radar-based CHMs (e.g., TanDEM-X) achieve global coverage, but are affected by canopy height saturation in dense forests and depend on high-quality terrain models to accurately isolate canopy height [131]. Optical data, while widely available, lack inherent sensitivity to vertical structure and require indirect proxies or fusion with information from active sensors  [158, 96] - yet none of the three most widely used global CHM models currently implement sensor fusion [84, 121, 148]. Most approaches instead rely on a single sensor type and derive height predictions either from temporal composites or from large numbers of hand-engineered features (e.g., >500>500> 500 features in Potapov et al. [121]), which may generalize poorly across ecosystems. These models tend to underestimate tall canopies, especially in tropical forests where most of the above-ground carbon is stored, and face challenges from persistent cloud cover, sensor saturation [83], and inconsistencies in input quality. As a result, significant discrepancies remain between global CHMs and airborne LiDAR benchmarks, and current products underestimate local variability and structural extremes [107]. Most CHMs are also static, failing to capture inter-annual dynamics, and inconsistencies in methodology and sensor characteristics hinder cross-product integration. These limitations underscore the need for scalable, temporally resolved, and sensor-fused CHMs with quantified uncertainties to support forest monitoring, carbon accounting, and Earth system modelling [108, 116].
@@
We evaluated TESSERA embeddings to predict canopy height at 10 m resolution within a 5 $\times$ 6 km area of tall old-growth tropical forests in the Danum Valley, Borneo [74] with openly available ground-truth data [33]. To ensure spatial independence, we conducted four-fold spatial cross-validation against the airborne LiDAR-derived canopy height, each time holding out a contiguous 50% of the region for testing and using the remaining 50% for training and validation. A 30-million-parameter UNet was trained using 32 $\times$ 32 pixel patches of the embeddings as input. To account for variability in training convergence, we performed three independent training runs per fold, each for 200 epochs, and retained the model with the best performance on the validation set. The accuracy metrics for the foundation models are reported as the average across the 12 runs. TESSERA achieved an $R^{2}$ of 0.66, a root mean squared error (RMSE) of 8.88 m, and a mean bias of 0.62-0.62- 0.62 m.
@@
We benchmarked TESSERA’s performance against state-of-the-art approaches through three sets of comparisons (Fig. 3). First, we evaluated its predictions against three global canopy height products: the GLAD map by Potapov et al. [121], the ETH map by Lang et al. [84], and the Meta map by Tolan et al. [148]. Although local models generally outperform global products in regional evaluations [127], the pronounced saturation of these global models at much lower canopy heights - combined with RMSEs that exceed their reported global averages [121, 84, 148]—underscores the challenge of accurately mapping structurally complex high canopy regions such as Danum Valley.
@@
We also compared our results to regional wall-to-wall canopy height maps for Indonesia, Malaysia, and the Philippines produced by Lang et al. [85], which used deep convolutional neural networks trained on Sentinel-2 time series and GEDI LiDAR to estimate regional canopy height and biomass. Second, we compared TESSERA to two alternative pixel-based foundation models, PRESTO [149] and Google Satellite Embedding (GSE) [56], using the same U-Net architecture, but replacing TESSERA’s input representations with those of the other foundation models. Finally, to assess the added value of the learned representations over the raw input data, we trained the same U-Net using cloud-free annual median composites of Sentinel-2 imagery.
@@
In all comparisons, TESSERA outperformed competing methods. Moreover, its embeddings do not require pre-processing, greatly enhancing reproducibility. Instead, only the region and year of interest are needed to replicate the input data pipeline. By fusing Sentinel-1 and Sentinel-2 data without requiring preprocessing or feature engineering, TESSERA’s 128-dimensional embeddings encode cloud-free, information-rich signals that effectively address key challenges in canopy height modelling.
@@
2.3 Above-ground biomass estimation
Forests play a central role in the global carbon cycle, creating a pressing need for reliable and scalable AGB estimates for climate modeling, carbon accounting, and informed strategies for conservation and sustainable land use [66]. Accurate estimation of above-ground biomass (AGB) is essential to quantify forest carbon stocks and monitor temporal changes due to deforestation, degradation, and recovery [45].
@@ 2.3 Above-ground biomass estimation
We evaluated the performance of TESSERA in predicting AGB against two leading models on a benchmark dataset. Firstly, we considered the geospatial foundation model SpectralGPT [67] combined with a UPerNet regression head – a setup found in the Pangaea benchmark [101] to outperform other foundation models tested on the BioMassters dataset [111]. As the preprocessing pipeline is not yet available in the codebase of Ref. [67] at the time of our analysis, we adopted the implementation from Ref. [101], which uses the preprocessed satellite imagery from the BioMassters dataset as input. Secondly, we used the winning model from the BioMassters competition as our state-of-the-art supervised baseline – a UNet with a temporal attention encoder [111], which achieved the lowest RMSE among over 1,000 submissions. The BioMassters dataset comprises 11,462 image patches that represent above-ground biomass in Finnish forests from 2017 to 2021. Ground-truth AGB values are derived from airborne LiDAR and aerial imagery, using calibrated allometric equations informed by extensive field plots by the Finnish Forest Centre and National Land Survey. Each patch covers a $2560\times 2560\,\textrm{m}^{2}$ area, with a spatial resolution of 10 m per pixel. The fine spatial granularity of the dataset makes the estimation task more challenging, as it introduces a long, thin tail of high AGB values that would not appear in coarser-resolution data.
A comparison of RMSE across varying training label fractions (Fig. 4a) shows that a UNet trained on the TESSERA embeddings consistently surpasses a UPerNet trained on the SpectralGPT representations, achieving lower prediction errors on 1% of the training set than SpectralGPT on the full training set. These results underscore the robustness of TESSERA in sparse-label regime. Moreover, TESSERA achieves results comparable to those of the supervised baseline trained on the full dataset, even with limited label availability. In general, bespoke models are inherently difficult to surpass in fully supervised settings. It is worth noting that the comparison to benchmarks is not entirely straightforward, as the satellite data used in the benchmark models differs from the TESSERA input. However, these discrepancies can be considered as an inherent part of the respective processing pipelines and hence largely unavoidable.
Figure 4: TESSERA embeddings demonstrate robustness for AGB mapping under limited label availability. a, TESSERA achieves lower RMSE values than the foundation model benchmark SpectralGPT. Following the evaluation protocol in Ref. [111], RMSE is computed on a fixed test set of 2773 patches and reported as the average RMSE per patch. For limited labels testing, a subset of labels is randomly chosen from the remaining data, and randomly split into training and validation sets (80:20). The figure shows the average RMSE computed over five random splits, with bars indicating the minimum–maximum range when it exceeds the size of the dot. The UNet model is trained on the TESSERA embeddings for 80 epochs, with the model achieving the lowest loss on the validation set selected as the final model. The inset depicts the coefficient of determination ($R^{2}$) and mean bias (MB) as functions of label availability. b, Predicted AGB map for an example patch using 10% of the training labels. c, The mean predicted AGB values show increasing bias at higher ground-truth AGB values, which are sparse in the dataset. The top panel depicts the histogram of the ground-truth AGB values. d, Pixel-wise comparison of predicted versus ground-truth AGB values for the test data with TESSERA+UNet (top), the supervised baseline (middle) and SpectralGPT+UPerNet (bottom).
@@ 2.4 Burned Area Detection
The frequency and severity of wildfires are projected to increase in many parts of the world due to climate change [5, 71, 117]. Accurate maps of burned areas and burn severity are critical for monitoring wildfire trends [10], predicting wildfire occurrence [29], assessing ecosystem recovery [186], and establishing effective fire management strategies [63]. Remote sensing is widely used to detect burned areas and assess severity regionally and globally [82].
@@
Most existing burn area mapping approaches rely on time series analysis or direct comparison between pre-fire and post-fire images, often requiring careful selection of suitable cloud-free images, manually tuned thresholds, and expert input [82, 119]. These methods typically use spectral indices such as the Normalized Burn Ratio (NBR), though many studies have also explored the design of custom features and applied spectral dimensionality reduction techniques to preserve key fire-related signals [82]. Most existing methods use optical sensors [82], which can detect fire-specific spectral signatures but are susceptible to cloud cover [64], particularly problematic for fire detection in ecosystems where vegetation rapidly re-greens [65]. SAR offers complementary sensitivity to vegetation structure and is unaffected by cloud cover [82, 143, 154]. Recent reviews emphasize the need for sensor fusion, combining passive and active sensors to improve accuracy and robustness [82, 30]. Sentinel-1 and 2 sensors in particular, when combined together at 10 m resolution, have shown promise in detecting burn areas from small fires that are often left out of coarse, global burn area products [143, 154, 129]. TESSERA extends the philosophy of traditional approaches by encoding temporal dynamics, while addressing major limitations through sensor fusion and the use of data-driven features learned directly from large volumes of combined Sentinel-1 and Sentinel-2 imagery (Fig. 5).
Figure 5: TESSERA embeddings capture wildfire extent, timing, and severity across two burned regions. (a–c) Burn Area A, which experienced two distinct fires in June and August 2021. (a) Time of fire map based on MTBS, showing unburned areas (blue), the June fire (orange), and the August fire (red). (b) Burn severity classification for the same region, with unburned (blue), low (green), moderate (orange), and high (red) severity. (c) UMAP projection of TESSERA embeddings from 2021, with colors corresponding to burn severity. Embeddings form well-separated clusters for unburned, June-, and August-burned areas, with a visible internal gradient reflecting severity within August fire. (d–f) Burn Area B, which experienced a single large fire in September 2021. (d) Time of fire map showing unburned (blue) and burned (green) areas. (e) Burn severity classification showing heterogeneous severity across the landscape. (f) UMAP projection for Burn Area B, revealing fuzzier boundaries between burned and unburned pixels, though the severity gradient remains evident.
We tested whether TESSERA’s temporal embeddings capture fire-induced changes in spectral and radar signals. We evaluated its ability to detect burned areas and burn severity using two randomly selected 15 $\times$ 15 km regions from the Monitoring Trends in Burn Severity (MTBS) dataset [4], each affected by wildfires in California during 2021. TESSERA embeddings for 2021 were obtained and UMAP dimensionality reduction was applied to a random subsample of pixels. Fire boundaries and burn severity data were sourced from the dataset. The resulting UMAP projections demonstrated strong separability in two-dimensional space, visually highlighting TESSERA’s capacity to detect disturbances at the pixel level. Burn Area A experienced two distinct wildfires—in June and August—while Burn Area B underwent a single large fire that affected mostly forested areas, resulting in a heterogeneous severity of the burn. The UMAP revealed separability across three domains (Fig. 5):
  • Burned area detection: In both regions, burned pixels formed distinct clusters from unburned, with a clearer separation in Burn Area A.
  • Temporal fire differentiation: In Burn Area A, embeddings separated pixels burned in June vs. August, suggesting temporal sensitivity to disturbance events.
  • Burn severity gradient: In both regions, a gradient structure in the UMAP space corresponded to burn severity, indicating that TESSERA encodes the magnitude of the disturbance.
Because TESSERA encodes an annual time series of spectral and radar signals for each pixel, it inherently captures changes in pixel trajectories without requiring prior preprocessing or the selection of cloud-free before-and-after images. This capacity supports a simplified, scalable pipeline for monitoring large-area disturbances.
@@ 2.5 Stocking indices in voluntary carbon markets
The conversion of degraded pastures into species-rich agroforestry systems has the potential to increase carbon sequestration while minimizing leakage liabilities [110]. To fund the transition, project developers can participate in the voluntary carbon market (VCM), but need to monitor above-ground biomass (AGB) over potentially large areas. This is a costly and inefficient undertaking if it is done only by field sampling.
To enable cost reductions, the most widely applied standard in the VCM (VM00047 from Verra [155]) allows project promoters to use remotely quantified ’stocking indices’ to assess project performance between 5-yearly ground truth assessments; project developers merely need to demonstrate that the chosen stocking index is correlated with the AGB. Upon acceptance, project developers can use the index to quantify the additionality of the project and to assess the performance of the project.
Here, we evaluate a simple TESSERA-derived stocking index against five widely used globally-available canopy height and AGB products that are often used in the VCM. To derive the stocking index from the TESSERA embeddings, GEDI relative height differences between RH95 and RH10 were used as target values for a pixel-wise RF regression following the methodology outlined in Atzberger et al. [13]. Without further fine-tuning, the derived stocking index, as well as five other state-of-the-art canopy height and AGB products (see  [1] for details), were individually rescaled so that their regression lines best matched (i.e., with slope 1 and zero bias) in situ data collected at 38 agroforestry sites in Para State, Brazil. This rescaling ensures a fair comparison between the various products.
@@
Figure 6: TESSERA outperforms various stocking indices commonly used in the Voluntary Carbon Market. Comparison of in-situ AGB measurements in 38 agro-forestry plots in Para State, Brazil with TESSERA (top left) and five other remote sensing products (see  [1] for details). Note that the different products were originally provided in different units and converted to AGB as follows: (a) TESSERA: difference in relative heights of GEDI estimates of RH90 and RH10 converted to AGB using 15.502 X + 160.5; (b) ETH canopy height converted using 3.4 X; (c) CTrees canopy height converted using 1.806 X + 44.9; (d) ESA AGB converted using 0.407 X + 34.0; (e) Meta canopy height converted using 3.723 X + 46.6; (f) GEDI-Tandem-X AGB prediction converted using 0.3 X + 54.3. Details on the conversion functions are provided in Supplemental section I.

### v3

@@ 2.2 Canopy height estimation
Despite substantial progress in global canopy height mapping, current approaches exhibit limitations that restrict their reliability and applicability. Airborne Laser Scanning (ALS), where laser-based LiDAR instruments are carried in an aircraft flying over a target region, creates highly accurate benchmark datasets, but is necessarily restricted in geographical range [23]. GEDI, a spaceborne LiDAR system, can provide accurate point measurements of canopy height to train models to map canopy height from optical and radar data; however, it does not provide wall-to-wall maps, does not sample high latitudes [42], and does not produce unreliable estimates in mountainous terrain [53]. Radar-based CHMs (e.g., TanDEM-X) achieve global coverage, but are affected by canopy height saturation in dense forests and depend on high-quality terrain models to accurately isolate canopy height [132]. Optical data, while widely available, may lack inherent sensitivity to vertical structure and require indirect proxies or fusion with information from active sensors  [160, 96] - yet none of the three most widely used global CHM models currently implement sensor fusion [84, 120, 150]. Most approaches instead rely on a single sensor type and derive height predictions either from temporal composites or from large numbers of hand-engineered features (e.g., >500>500> 500 features in Potapov et al. [120]), which may generalize poorly across ecosystems. These models tend to underestimate tall canopies, especially in tropical forests where most of the above-ground carbon is stored, and face challenges from persistent cloud cover, sensor saturation [83], and inconsistencies in input quality. As a result, significant discrepancies remain between global CHMs and ALS benchmarks, and current products underestimate local variability and structural extremes [106]. Most CHMs are also static, failing to capture inter-annual dynamics, and inconsistencies in methodology and sensor characteristics hinder cross-product integration. These limitations underscore the need for scalable, temporally resolved, and sensor-fused CHMs with quantified uncertainties to support forest monitoring, carbon accounting, and Earth system modelling [107, 115].
@@
We evaluated TESSERA embeddings to predict canopy height at 10 m resolution within a 5 $\times$ 6 km area of tall old-growth tropical forests in the Danum Valley, Borneo [75] with openly available ground-truth data [34]. To ensure spatial independence, we conducted four-fold spatial cross-validation against the airborne LiDAR-derived canopy height, each time holding out a contiguous 50% of the region for testing and using the remaining 50% for training and validation. A 30-million-parameter UNet was trained using 32 $\times$ 32 pixel patches of the embeddings as input. To account for variability in training convergence, we performed three independent training runs per fold, each for 200 epochs, and retained the model with the best performance on the validation set. The accuracy metrics for the foundation models are reported as the average across the 12 runs. TESSERA achieved a coefficiento of determination ($R^{2}$) of 0.66, a root mean squared error (RMSE) of 8.88 m, and a mean bias of 0.62-0.62- 0.62 m.
@@
We benchmarked TESSERA’s performance against state-of-the-art approaches through three sets of comparisons (Fig. 4). First, we evaluated its predictions against three global canopy height products: the GLAD map by Potapov et al. [120], the ETH map by Lang et al. [84], and the Meta map by Tolan et al. [150]. Although local regression models, such as ours, generally outperform global products in regional evaluations [128], the pronounced saturation of these global models at much lower canopy heights–combined with RMSEs that exceed their reported global averages [120, 84, 150]—underscores the challenge of accurately mapping structurally complex high canopy regions such as Danum Valley.
@@
We also compared our results with regional wall-to-wall canopy height maps for Indonesia, Malaysia, and the Philippines produced by Lang et al. [85], which used deep convolutional neural networks trained in Sentinel-2 time series and GEDI LiDAR to estimate the regional canopy height and biomass. Second, we compared TESSERA to two alternative pixel-based foundation models, PRESTO [151] and Google Satellite Embedding (GSE) [57], using the same U-Net architecture, but replacing TESSERA’s input representations with those of the other foundation models. Finally, to assess the added value of the learned representations over the raw input data, we trained the same U-Net using cloud-free annual median composites of Sentinel-2 imagery.
@@
In all comparisons, TESSERA outperformed competing methods. Moreover, its embeddings do not require pre-processing, greatly enhancing reproducibility. Instead, only the region and year of interest are needed to replicate the input data pipeline. By fusing Sentinel-1 and Sentinel-2 data without requiring preprocessing or careful selection of input features, TESSERA’s 128-dimensional embeddings encode cloud-free, information-rich signals that effectively address key challenges in canopy height modelling.
@@
2.3 Above-ground forest biomass estimation
Forests play a central role in the global carbon cycle [67]. This motivates the need for reliable and scalable above-ground biomass (AGB) estimates of forest carbon stocks to allow monitoring of deforestation, forest degradation, and forest recovery [46]. A critical problem in AGB estimation is that, as with canopy height estimation, precise values are best obtained from field plots or airborne laser scanning, which are too costly to scale beyond small geographic regions. Thus, there has been a long-standing effort to link field measurements to global satellite observations [194].
Recently, with the creation of the BioMassters benchmark dataset [110] (more on this below), a global challenge was issued to research groups to predict AGB in Finnish forest patches from Sentinel-1 and Sentinel-2 satellite observations using a machine learning approach of their choice  [5]. We found that the winner, chosen from more than 1,000 submissions, achieved a mean RMSE value of 25.93 tonnes/hectare (t/ha), a substantial improvement over the prior benchmark score of 101.98. In our work, we compared a regression model trained on TESSERA embeddings with this winning model, as well as regression models trained on embeddings from two other foundation models: SpectralGPT [68] and GSE [57].
@@ 2.3 Above-ground forest biomass estimation
First, a summary of the BioMassters dataset. This data set contains 11,462 patches of Finnish forests from 2017 to 2021, with associated ground truth AGB values provided by the Finnish Forest Center and the National Land Survey derived from airborne LiDAR scanning (ALS), aerial imagery and calibrated allometric equations [126, 127] informed by extensive field plots. Each patch covers a $2560\times 2560\,\textrm{m}^{2}$ area, with a spatial resolution of 10 m per pixel. The fine spatial granularity of the dataset makes the estimation task more challenging, as it introduces a long, thin tail of high AGB values (Fig. 5b) that would not appear in coarser resolution data. To reduce the influence of implausibly high AGB values, we clip values above 500 t/ha to this threshold.
We compared a UNet-based regression model [129] trained on TESSERA embeddings with SpectralGPT embeddings [68] combined with a lightweight spatial-temporal encoder [133] for temporal aggregation and a UPerNet [181] regression head, because this combination has been found in the Pangaea benchmark [100] to outperform other foundation models [110, 100]. Unfortunately, as of the time of this writing, SpectralGPT’s preprocessing pipeline has not yet been published in their codebase. Thus, we adopted the Pangaea implementation of their model [100], which uses preprocessed satellite imagery from the BioMassters dataset. We also compared TESSERA with the model that won the BioMassters competition as our state-of-the-art supervised baseline. This model uses a UNet with a temporal attention encoder [110]. Our comparison of TESSERA to the supervised baseline and SpectralGPT is not entirely straightforward, as the snapshot satellite data used in these models differs from the annual time series required as TESSERA’s input. However, these discrepancies can be considered as an inherent part of the respective processing pipelines and hence largely unavoidable. Finally, we compared our work with a UNet-based regressor trained on GSE embeddings.
Our performance metrics are the mean RMSE, the coefficient of determination ($R^{2}$), and the mean bias (MB) in AGB across the test patches as a function of the fraction of labelled data used to train the regressor (Fig. 5a). We make three significant findings. First, although custom models developed using supervised learning are typically difficult to outperform, even with a small fraction of training labels, our regressor nearly matches the performance of the BioMassters competition winner, which uses a supervised learning approach. Second, our regressor trained on 1% of the labelled data outperforms a UPerNet-based regressor trained on the entire training set using SpectralGPT embeddings. Third, our regressor achieves performance comparable to GSE, a commercially developed foundation model with far more training resources at its disposal.
Figure 5: A regressor trained on TESSERA embeddings matches the performance of winner of the BioMassters competition with far fewer labels, outperforms SpectralGPT+UPerNet, the best performing foundation model in the BioMassters challenge, and matches the performance of the commercially-developed GSE embeddings. a, TESSERA achieves lower RMSE values than SpectralGPT+UPerNet and results similar to the state-of-the-art supervised baseline and Google Satellite Embedding foundation model (top row). The coefficient of determination ($R^{2}$) and mean bias (MB) are shown in the middle and bottom row as functions of label availability. b, Mean predicted AGB values show increasing bias at higher ground-truth AGB values, which, however, are sparse in the dataset (the y-axis in the upper panel is on a logarithmic scale). Mean predictions are calculated for ground-truth AGB values grouped into 5 t/ha intervals. c, Predicted AGB map for an example patch using 4% of labels in training. d, Pixel-wise comparison of predicted versus ground-truth AGB values for the test data with TESSERA+UNet (left) and the supervised baseline (right). The supervised baseline ignores ground-truth values above 400 t/ha during training, which is reflected in its lower maximum predictions.
@@ 2.4 Burned Area Detection
The frequency and severity of wildfires are projected to increase in many parts of the world due to climate change [6, 72, 116]. Accurate maps of burned areas and burn severity are critical for monitoring wildfire trends [11], predicting wildfire occurrence [30], assessing ecosystem recovery [188], and establishing effective fire management strategies [64].
@@
Remote sensing is widely used to detect burned areas and assess severity regionally and globally [82]. Most existing burn area mapping approaches rely on time-series analysis or direct comparison between pre-fire and post-fire images, often requiring careful selection of suitable cloud-free images, manually tuned thresholds, and expert input [82, 118]. These methods typically use spectral indices such as the Normalized Burn Ratio (NBR), although many studies have also explored the design of custom features and applied spectral dimensionality reduction techniques to preserve key fire-related signals [82]. Most existing methods use optical sensors [82], which can detect fire-specific spectral signatures but are susceptible to cloud cover [65], particularly problematic for fire detection in ecosystems where vegetation rapidly re-greens [66]. SAR offers complementary sensitivity to vegetation structure and is unaffected by cloud cover [82, 145, 156]. Recent reviews emphasize the need for multimodal remote sensing, combining passive and active sensors to improve accuracy and robustness [82, 31]. Sentinel-1 and -2 sensors in particular, when combined together at 10 m resolution, have shown promise in detecting burn areas from small fires that are often left out of coarse, global burn area products [145, 156, 130]. TESSERA extends these approaches by encoding temporal dynamics, while addressing major limitations through the use of multimodal sensors and the use of data-driven features learned directly from large volumes of combined Sentinel-1 and Sentinel-2 time series (Fig. 6).
Figure 6: Fire disturbance detection using TESSERA embeddings across two burned regions in California. a, UMAP projection of 2020 TESSERA embeddings reveals clear structure separating unburned pixels and two temporally distinct fire events. A continuous gradient within the August fire aligns with MTBS burn severity classes, suggesting that TESSERA captures both timing and magnitude of disturbance. b, Burned vs. unburned classification performance using random forest classifiers trained on different input representations. TESSERA achieves high F1-scores with minimal training data and performs comparably to a task-specific composite baseline, while significantly outperforming Google Satellite Embeddings (GSE). c, Burn classification maps generated from models trained on 1% of labelled data demonstrate strong agreement with MTBS perimeters across both test areas. As MTBS is itself derived from remote sensing imagery, discrepancies may reflect methodological differences rather than prediction error.
We tested whether TESSERA embeddings capture fire-induced disturbance by evaluating their ability to detect burned areas and severity of the burn (Figure 6). Our analysis focused on two randomly selected 15$\times$15 km regions in California that experienced wildfires in 2020, using fire boundaries and burn severity labels from the Monitoring Trends in Burn Severity (MTBS) dataset [4]. UMAP projections [102] of the TESSERA embeddings revealed a strong separability between burned and unburned pixels, with individual fires forming distinct clusters. In particular, a gradient in the UMAP space was aligned with the severity of the burn in both areas, indicating that TESSERA encodes not only the occurrence but also the magnitude of the disturbance (Figure 6a). To quantify separability, we trained a Random Forest classifier to distinguish burned pixels from unburned pixels. We compared three input representations: (1) TESSERA embeddings, (2) Google Satellite Embeddings (GSE), and (3) a task-specific baseline constructed from pre- and post-fire Sentinel-2 composites. The classifier trained in TESSERA embeddings achieved an average F1 score of more than 0.96 (Figure 6 b). Remarkably, it reached over 90% of its final performance using only 0.01% of the training labels ( 330 pixels), demonstrating strong data efficiency. These results show that TESSERA’s annual task-agnostic embeddings perform on par with a task-specific baseline that requires prior knowledge of fire timing and spectral properties, while significantly outperforming other commercially developed foundation models such as Google Satellite Embeddings.
To assess practical utility, we used models trained in 1% of the data to generate wall-to-wall burn maps. These predictions closely matched MTBS perimeters and exhibited minimal false positives in the unburned areas (Figure 6C). The results confirm that TESSERA’s general-purpose representations encode high-resolution disturbance signatures and support efficient, scalable fire mapping workflows without relying on handcrafted features or task-specific adjustments.
@@ 2.5 Stocking indices in voluntary carbon markets
The conversion of degraded pastures into species-rich agroforestry systems has the potential to increase carbon sequestration without displacing cattle grazing activities elsewhere [109]. To fund the transition, project developers can participate in the voluntary carbon market (VCM), but need to monitor above-ground biomass (AGB) over potentially large areas. This is a costly and inefficient undertaking if it is done only by field sampling.
To enable cost reductions, the most widely applied standard in the VCM (VM00047 from Verra [157]) allows project promoters to use remotely quantified ’stocking indices’ to assess project performance between 5-yearly ground truth assessments; project developers need to demonstrate that the chosen stocking index is correlated with the AGB. Upon acceptance, project developers can use the index to assess the performance of the project. Here, we evaluate a simple TESSERA-derived stocking index against five widely used globally-available canopy height and AGB products that are often used in the VCM. To derive the stocking index from the TESSERA embeddings, GEDI relative height differences between RH95 and RH10 were used as target values for a pixel-wise random forest regression following the methodology outlined in Atzberger et al. [14]. Without further fine-tuning, the derived stocking index, as well as five other state-of-the-art canopy height and AGB products (see  [1] for details), were individually rescaled so that their regression lines best matched (i.e., with slope 1 and zero bias) in situ data collected at 38 agroforestry sites in Para State, Brazil. This rescaling ensures a fair comparison between the various products.
@@
Figure 7: TESSERA outperforms various stocking indices commonly used in the Voluntary Carbon Market. Comparison of in-situ AGB measurements in 38 agro-forestry sites, within which 1-3 inventory plots were measured in Para State, Brazil with TESSERA (top left) and five other remote sensing products (see  [1] for details). Note that the different products were originally provided in different units and converted to AGB as follows: (a) TESSERA: difference in relative heights of GEDI estimates of RH90 and RH10 converted to AGB using 15.502 X + 160.5; (b) ETH canopy height converted using 3.4 X; (c) CTrees canopy height converted using 1.806 X + 44.9; (d) ESA AGB converted using 0.407 X + 34.0; (e) Meta canopy height converted using 3.723 X + 46.6; (f) GEDI-Tandem-X AGB prediction converted using 0.3 X + 54.3. Details on the conversion functions are provided in Supplemental section I.